Charlies Furniture Store has been in business for several years. The firms owners have described the store as a high-priceI only need help with part C and D
Answer #1

a. Turnover is @ 0.6 of assets. So Turnover => 800,000*0.6 = 480,000. Margin => 480000*29% = 139200. SO, ROI = 139200/800,000*100 = 17.4%

b. Since ROI and Total average assets remain the same, Required margin = 139200. This is the margin ie., 20%. So sales => 139200/20*100 = 696000.

c. Total Increase in sales amount => 696000 (ie b) - 480000 (ie a) = 216000 $. In % of existing sales => 216000/450000*100 = 45%. So the increase is of 45% in sales and not 50% as Charile said.

d. Advertising extensively and incurring market research expenses will increase the costs. Thus the incremental income generated should first recover the costs incurred on product and then provide for the costs incurred in advertising and research. Moreover, it has to provide the margin required and the required ROI.